OpenStudy (anonymous):

Slove by the substution method 2m+n=5 m-5n=30

OpenStudy (anonymous):

m = 5 n = -5

OpenStudy (cathyangs):

2m+n=5 n=5-2m m-5(5-2m)=30 m-25+10m=30 11m=55 m=5 2(5)+n=5 10+n=5 n=-5

OpenStudy (diyadiya):

m=30+5n substitute it in 1st equation 2(30+5n)+n=5 60+10n+n=5 11n=-55 n=-5 2m-5=5 2m=10 m=5